Welcome to the CEDA news channel! We will post service announcements, highlights about our work, training events, vacancies and other important things here.
Temporary BADC file server shutdown on Tuesday 12th December from 5pm ====================================================================== The BADC File server (tornado) will be temporarily shutdown on ** Tuesday 11th December 2001 from 5pm. ** This temporary interruption is necessary for some hardware upgrade. ** Normal service should resume on Wednesday 12th December 2001 ** This shutdown unfortunately implies that the BADC Web site and access to data and services will be temporarily unavailable. We are sorry for the inconvenience this may cause. BADC Support 07/12/01.Read more
Northern Hemisphere Mean Sea Level Pressure data now available ============================================================== This data set holds gridded (5° latitude by 10° longitude grid) Northern Hemisphere (north of 15° N) monthly and daily series of Mean Sea Level Pressure (MSLP) fields. The monthly series data are available for the period 1873 to 1997; the daily series data are available for the period 1881 to 1997 . More information available at http://www.badc.rl.ac.uk/data/mslp/ BADC Team 04/12/2001Read more
Historical Northern Hemisphere Geopotential Height data now available from 1945 =============================================================================== This data set holds gridded (5 degree latitude by 5 degree longitude grid) Northern Hemisphere (north of 15N) daily and monthly series of 500hPa geopotential height fields and also daily and monthly series of 1000-500hPa thickness fields. The historical data is now available for the period 1945 to present. More information available at http://www.badc.rl.ac.uk/data/height/ BADC TeamRead more
Dear BADC User, This message is to inform you that the BADC File server (tornado) will be temporarily shutdown from Friday 16th November at 3pm to Sunday 18th November. The BADC file server is expected to be operational again from Monday 19th November. This shutdown unfortunately implies that the BADC Web site and access to data and services will be temporarily unavailable. We are sorry for the inconvenience this may cause. Regards, BADC Support BADC Support 15/11/01.Read more
Message forwarded from Dr. Charles H. Jackman, NASA/Goddard Space Flight Center. ================================================================================ October 25, 2001 UARS Colleagues: We have generally positive news to report. Please send this on to others as I may not have everyone on my UARS distribution list. The latest news regarding the UARS Mission is given below: MEETING AT NASA HEADQUARTERS ON OCTOBER 24TH - We had a meeting yesterday at NASA Headquarters with Mike Luther (Deputy Associate Administrator) and others in Code Y. We discussed the status of the UARS and the many changes that are ongoing in the project as well as this option. Generally we received positive feedback. We have been told to proceed with the "Traceability Mission" option (see below) for the foreseeable future. The total cost of the "Traceability Mission" option including mission and science operations is about $3.7M. For comparison, the cost of the UARS mission and science operations was about $15.4M in FY01. We have received $2M so far that is being applied to the mission operations. UARS "TRACEABILITY MISSION" OPTION - This option will be primarily focussed on providing validation/calibration measurements for upcoming future missions. We hope to provide UARS validation measurements for the following future FY02 missions: TIMED, SAGE III, ENVISAT-1, ADEOS II, EOS Aqua, and SORCE. If we are successful in FY02, it may be possible (with Code Y approval) to continue for an additional time period beyond September 30, 2002, and overlap other missions. UARS INSTRUMENT DATA COLLECTION - Instrument science observations resumed on October 1st after being shutdown from September 24-30. We are now in a "baseline operational mode" wherein we have 15 contacts/day (once per orbit) as compared with 30 contacts/day (twice per orbit) in FY01. This allows the collection of about 20-30% of the possible data. Since the HALOE instrument uses the remaining working part of the tape recorder, practically all of the HALOE observations are being collected. There is enough power to operate HALOE on about 2/3rds of the days. After the launches of new missions, with an appropriate time for outgassing and new instrument checks, the UARS instruments will be operated in an "intense validation/calibrational mode" for two-week observational periods. During these periods, the UARS instruments will have 30 contacts/day so that 50-55% of the data would be collected. UARS CENTRAL DATA HANDLING FACILITY (CDHF) - Unfortunately, we will not be able to fund the UARS CDHF beyond November 30, 2001. The UARS CDHF is in the midst of final reprocessing, data archive, and supporting the effort to move to data distribution through a "new" system called PACOR-A. UARS LONG-TERM DATA ARCHIVE - This issue is currently being addressed. It is hoped that ultimately all the UARS data (including L0/L1/L2) will be moved to the DAAC and be compatible with UNIX machines. To complete this effort will require some funding beyond what has been approved. UARS RETRIEVAL - It was decided at the meeting yesterday that there will not be a Space Shuttle retrieval of the UARS. Thanks for your support. I am convinced that the community efforts over these difficult several months. Sincerely, Charles Jackman UARS Project ScientistRead more